OpenStudy (anonymous):

x = -135, cos 2x = 1 - 2 sin2x True False I think its true, but when I was using the unit circle, I couldnt find -135

OpenStudy (anonymous):

@Chlorophyll can you help me, please?

OpenStudy (anonymous):

hi friend, your problem is \[\cos 2x = 1 - 2 \sin^2 x ?\]

OpenStudy (anonymous):

if mine is right, so, the answer is true. if yours is right, the answer is false. why? because we have a formula for that trig: it's mine. it's true with any value of x. Hope this helps

OpenStudy (anonymous):

Thanks you, I was having a big problem with this, and I moved to the next questions. Thank YOU SO MUCH!

OpenStudy (anonymous):

thanks for your medal.